# Holographic thermodynamics of a five-dimensional neutral Gauss-Bonnet AdS black hole

### v1

@@ 3.1 Canonical ensemble: swallowtail behavior and critical point
$\mathcal{C}_{C,\mathcal{V},\tilde{\alpha}}=T\left(\frac{\partial S}{\partial\tilde{T}}\right)_{C,\mathcal{V},\tilde{\alpha}}=\frac{12\pi\text{C}\left(2x^{3}+x\right)\left(x^{2}+2y\right)^{2}}{x^{2}(12y-1)+2x^{4}+2y}.$ (31)
The heat capacities for the low and high entropy branches of the Helmholtz free energy in Fig. 1 are both positive, hence they are thermodynamically stable. While the heat capacity for the intermediate entropy branch is negative, hence it is thermodynamically unstable. We also draw the heat capacity for the CFT states as in Fig. 2. For the brown dashed and blue line, the central charge $C$ is larger than the critical value $C_{\text{c}}=9\pi\tilde{\alpha}/\left(2\sqrt{G_{\text{N}}}\right)$. For the red large dashing line, the central charge is equal to the critical value $C=C_{\text{c}}$. For the black dotted line, the central charge $C$ is smaller than the critical value $C_{\text{}c}$. As we can see from the figure, the heat capacity is continues and finite for small central charge $C<C_{\text{c}}$. However, there are points where the heat capacity is divergent for large central charge $C\geq C_{\text{c}}$. The divergent points of the heat capacity correspond to the intersection points between the small/large entropy branches and intermediate entropy branches in Fig. 1.
Figure 2: The heat capacity for the thermodynamic states of the boundary CFT, where we have set the Newtown’s constant $G_{\text{N}}=1$, the boundary radius $R=10,000$ and the CFT thermodynamic variable $\tilde{\alpha}=1$. For different curves the central charge are different. The heat capacity is continues and finite for small central charge $C<C_{\text{c}}$; while for large central charge $C>C_{\text{c}}$, the heat capacity consists of three parts and there are points where the heat capacity is divergent.
@@
$\begin{split}\tilde{T}_{\text{c}}&=\frac{\sqrt{6}}{2\pi R},\qquad\quad\tilde{\mathcal{A}}_{\text{c}}=-\frac{9\pi}{8\sqrt{G_{\text{N}}}R},\qquad\mathcal{P}_{\text{c}}=\frac{\tilde{\alpha}}{2\pi\sqrt{G_{\text{N}}}R^{4}},\\ S_{\text{c}}&=\frac{\sqrt{6}}{\sqrt{G_{\text{N}}}}\pi^{2}\tilde{\alpha},\qquad C_{\text{c}}=\frac{9\pi}{2\sqrt{G_{\text{N}}}}\tilde{\alpha},\quad\quad\mu_{\text{c}}=\frac{1}{4R}.\end{split}$ (35)
@@
3.2 Maxwell’s equal area law and coexistence curve
The physical processes of crossing the coexistence phase often unveil intriguing phenomena, such as latent heat Wei:2015iwa ; Wei:2023mxw . As shown by previous work Zhou:2020vzf ; Wei:2014qwa ; Zhou:2019xai , both the swallowtail behavior of the free energy and Maxwell’s equal area law can reveal phase transition of a thermodynamic system and they are consistent with each other. In this subsection, we derive Maxwell’s equal area law from the change in the Helmholtz free energy during the phase transition between low and high entropy states. Then we employ Maxwell’s equal area law to construct an analytical formula for the coexistence curve in the $\tilde{T}-S$ plane.
@@ 3.2 Maxwell’s equal area law and coexistence curve
$\begin{split}\tilde{T}(\text{A})\left[S(\text{C})-S(\text{A})\right]=\frac{2C}{R}\frac{2x_{\text{L}}^{3}+x_{\text{L}}}{x_{\text{L}}^{2}+2y}(x_{\text{L}}^{2}+x_{\text{L}}x_{\text{H}}+x_{\text{H}}^{2}+6y)(x_{\text{H}}-x_{\text{L}}).\end{split}$ (44)
@@
$\begin{split}3(x_{\text{L}}+x_{\text{H}})(x_{\text{L}}^{2}+2y)(x_{\text{L}}^{2}+x_{\text{H}}^{2}+1)=2(2x_{\text{L}}^{3}+x_{\text{L}})(x_{\text{L}}^{2}+x_{\text{L}}x_{\text{H}}+x_{\text{H}}^{2}+6y).\end{split}$ (45)
@@
$\begin{split}\tilde{T}=\frac{\sqrt{1-24y-\sqrt{48y(9y-1)+1}}\left(2-24y-\sqrt{48y(9y-1)+1}\right)}{\sqrt{2}\pi R\left(1-20y-\sqrt{48y(9y-1)+1}\right)}.\end{split}$ (48)
@@
$\tau=\frac{\sqrt{-\sqrt{3}\sqrt{\frac{1}{c^{2}}-\frac{4}{c}+3}-\frac{2}{c}+3}\left(c\left(\sqrt{3}\sqrt{\frac{1}{c^{2}}-\frac{4}{c}+3}-6\right)+2\right)}{c\left(3\sqrt{3}\sqrt{\frac{1}{c^{2}}-\frac{4}{c}+3}-9\right)+5}.$ (51)
One of the most remarkable properties of the above equation for the coexistence curve is its independence of the CFT thermodynamic variable $\tilde{\alpha}$, as well as the boundary volume $\mathcal{V}$. From the coexistence curve in Fig. 5, the temperature for the low and high entropy coexistence phases decreases with the increasing of the inverse of the central charge $1/C$.
@@ 3.3 Critical exponents
$S_{\text{L}}$ $=\sqrt{2}\pi C\sqrt{1-24y-\sqrt{48y(9y-1)+1}}\left(1-12y-\sqrt{48y(9y-1)+1}\right),$ (57)
@@
$\begin{split}\eta&=\frac{S_{\text{H}}-S_{\text{L}}}{S_{\text{c}}}\\ &=\frac{\left(12y+\sqrt{48y(9y-1)+1}-1\right)^{2}\left(36y+\sqrt{48y(9y-1)+1}-1\right)}{8\sqrt{3}y\left(-24y-\sqrt{48y(9y-1)+1}+1\right)^{3/2}}.\end{split}$ (59)
@@
$\begin{split}t&=\frac{\tilde{T}}{\tilde{T}_{\text{c}}}-1\\ &=\frac{\sqrt{1-24y-\sqrt{48y(9y-1)+1}}\left(24y+\sqrt{48y(9y-1)+1}-2\right)}{\sqrt{3}\left(20y+\sqrt{48y(9y-1)+1}-1\right)}-1.\end{split}$ (60)
@@
$\begin{split}\eta&=\frac{4\left(t(t+2)-\sqrt{3}\sqrt{t(t+1)^{2}(t+2)}+1\right)^{2}\left(\sqrt{3}\sqrt{t(t+1)^{2}(t+2)}-3t(t+2)\right)}{(1-2t(t+2))\left(4t(t+2)-2\sqrt{3}\sqrt{t(t+1)^{2}(t+2)}+1\right)^{3/2}}\\ &=4\sqrt{6}t^{\frac{1}{2}}+25\sqrt{6}t^{\frac{3}{2}}+\mathcal{O}(t^{\frac{5}{2}}).\end{split}$ (62)
@@
$\begin{split}\chi\thickapprox\frac{(1+3c)(c-1)^{\frac{1}{3}}}{3c^{\frac{4}{3}}}\sim\left(c-1\right)^{\frac{1}{3}},\end{split}$ (63)
@@
Thus, we have obtained the critical exponents $\alpha^{\prime}$, $\beta$ and $\delta$ associated with the $\tilde{T}-S$ criticality of the CFT, where the conformal factor $\omega$ is free to vary but with fixed Newton’s constant. Our results suggest that these critical exponents associated with the $\tilde{T}-S$ criticality of the CFT are the same as those of the five-dimensional Gauss-Bonnet AdS black hole CCLY13 ; Miao:2018fke .
4 Conclusion and discussion
In the AdS/CFT correspondence, black holes in AdS spacetime correspond to thermal states of the boundary field theories. Based on the recently proposed viewpoint of the holographic thermodynamics, which treats the conformal factor $\omega$ as a variable, we derived the first law and Euler equation for the thermodynamics of the CFT, and established a holographic dictionary between the thermodynamics of the five-dimensional neutral Gauss-Bonnet AdS black hole and the dual boundary thermodynamics. Having the thermodynamics for the dual CFT, we investigated the phase transition and criticality of the CFT in the fixed $(C,\mathcal{V},\tilde{\alpha})$ canonical ensemble. Just like the free energy for the five-dimensional neutral Gauss-Bonnet AdS black hole, the free energy for the CFT displays a characteristic swallowtail behavior. This indicates a first-order phase transition below the critical point and a second-order phase transition at the critical point. Through the $\tilde{T}-S$ oscillatory behavior, we obtained the critical point for the thermodynamics of the CFT in the $\tilde{T}-S$ plane. The critical point for the CFT is exact the same with that of the five-dimensional neutral Gauss-Bonnet AdS black hole. Using Maxwell’s equal area law, we got the coexistence curve for the high and low entropy phases of the CFT. Besides, we got the critical exponents for the CFT, and found that these critical exponents associated with the $\tilde{T}-S$ criticality are the same as those of the five-dimensional Gauss-Bonnet AdS black hole.
Critical exponents describe the singularity behavior of thermodynamic variables such as the heat capacity, order parameter etc. They are universal and satisfy the scaling hypothesis. The critical exponent $\gamma$ governing the isothermal compressibility of black hole thermodynamics is defined as
@@ 4 Conclusion and discussion
As we argued in Sec. 2.2, the black hole thermodynamic pressure $P$ is dual to the CFT central charge $C$, and the black hole thermodynamic volume $V$ is dual to the CFT chemical potential $\mu$. In our investigation of the critical exponents for the CFT thermodynamics in the $\tilde{T}-S$ plane, we found that if we define the critical exponent $\gamma$ just replacing the pressure $P$ and the specific volume $v$ by the central charge $C$ and the chemical potential $\mu$ respectively
@@
Criticality is a very important phenomena for thermodynamic systems and critical exponents are universal. Our investigation of the thermodynamics and criticality from the gravity/gauge duality viewpoint might lead to new insight to the understanding of the Van der Waals like phase behavior of the five-dimensional Gauss-Bonnet AdS black hole.
@@ Acknowledgements.
We are grateful to Chuan Chen, Shan-Ping Wu, and Yu-Peng Zhang for useful discussions. This work was supported by National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12305065, No. 12247178, No. 12347177, No. 12075103 and No. 12247101), the China Postdoctoral Science Foundation (Grant No. 2023M731468), the 111 Project (Grant No. B20063), and Lanzhou City’s scientific research funding subsidy to Lanzhou University.

### v2

@@ III.1 Canonical ensemble: swallowtail behavior and critical point
$\begin{split}\mathcal{C}_{C,\mathcal{V},\tilde{\alpha}}&=T\left(\frac{\partial S}{\partial\tilde{T}}\right)_{C,\mathcal{V},\tilde{\alpha}}\\ &=\frac{12\pi\text{C}\left(2x^{3}+x\right)\left(x^{2}+2y\right)^{2}}{x^{2}(12y-1)+2x^{4}+2y}.\end{split}$ (32)
The heat capacities for the low and high entropy branches of the Helmholtz free energy in Fig. 1 are both positive, hence they are thermodynamically stable. While the heat capacity for the intermediate entropy branch is negative, hence it is thermodynamically unstable. We also draw the heat capacity for the CFT states as in Fig. 2. For the brown dashed and blue line, the central charge $C$ is larger than the critical value $C_{\text{c}}=9\pi\tilde{\alpha}/\left(2\sqrt{G_{\text{N}}}\right)$. For the red large dashing line, the central charge is equal to the critical value $C=C_{\text{c}}$. For the black dotted line, the central charge $C$ is smaller than the critical value $C_{\text{}c}$. As we can see from the figure, the heat capacity is continuous and finite for small central charge $C<C_{\text{c}}$. However, there are points where the heat capacity is divergent for large central charge $C\geq C_{\text{c}}$. The divergent points of the heat capacity correspond to the intersection points between the small/large entropy branches and intermediate entropy branches in Fig. 1.
Figure 2: The heat capacity for the thermodynamic states of the boundary CFT, where we have set the Newtown’s constant $G_{\text{N}}=1$, the boundary radius $R=10,000$ and the CFT thermodynamic variable $\tilde{\alpha}=1$. For different curves the central charge are different. The heat capacity is continuous and finite for small central charge $C<C_{\text{c}}$; while for large central charge $C>C_{\text{c}}$, the heat capacity consists of three parts and there are points where the heat capacity is divergent.
@@
$\begin{split}\tilde{T}_{\text{c}}&=\frac{\sqrt{6}}{2\pi R},\qquad\qquad\tilde{\mathcal{A}}_{\text{c}}=-\frac{9\pi}{8\sqrt{G_{\text{N}}}R},\\ \mathcal{P}_{\text{c}}&=\frac{\tilde{\alpha}}{2\pi\sqrt{G_{\text{N}}}R^{4}},\qquad S_{\text{c}}=\frac{\sqrt{6}}{\sqrt{G_{\text{N}}}}\pi^{2}\tilde{\alpha},\\ C_{\text{c}}&=\frac{9\pi}{2\sqrt{G_{\text{N}}}}\tilde{\alpha},\qquad\quad\mu_{\text{c}}=\frac{1}{4R}.\end{split}$ (36)
@@
III.2 Maxwell’s equal area law and coexistence curve
The physical processes of crossing the coexistence phase often unveil intriguing phenomena, such as latent heat [59, 60]. As shown by previous work [61, 62, 63], both the swallowtail behavior of the free energy and Maxwell’s equal area law can reveal phase transition of a thermodynamic system and they are consistent with each other. In this subsection, we derive Maxwell’s equal area law from the change in the Helmholtz free energy during the phase transition between low and high entropy states. Then we employ Maxwell’s equal area law to construct an analytical formula for the coexistence curve in the $\tilde{T}-S$ plane.
@@ III.2 Maxwell’s equal area law and coexistence curve
$\tilde{T}(\text{A})\left[S(\text{C})-S(\text{A})\right]=\frac{2C}{R}\frac{2x_{\text{L}}^{3}+x_{\text{L}}}{x_{\text{L}}^{2}+2y}\\ \times(x_{\text{L}}^{2}+x_{\text{L}}x_{\text{H}}+x_{\text{H}}^{2}+6y)(x_{\text{H}}-x_{\text{L}}).$ (45)
@@
$3(x_{\text{L}}+x_{\text{H}})(x_{\text{L}}^{2}+2y)(x_{\text{L}}^{2}+x_{\text{H}}^{2}+1)=2(2x_{\text{L}}^{3}+x_{\text{L}})\\ \times(x_{\text{L}}^{2}+x_{\text{L}}x_{\text{H}}+x_{\text{H}}^{2}+6y).$ (46)
@@
$\tilde{T}=\frac{\sqrt{1-24y-\sqrt{48y(9y-1)+1}}}{\sqrt{2}\pi R\left(1-20y-\sqrt{48y(9y-1)+1}\right)}\\ \times\left(2-24y-\sqrt{48y(9y-1)+1}\right).$ (49)
@@
$\tau=\sqrt{-\sqrt{3}\sqrt{\frac{1}{c^{2}}-\frac{4}{c}+3}-\frac{2}{c}+3}\\ \times\frac{c\left(\sqrt{3}\sqrt{\frac{1}{c^{2}}-\frac{4}{c}+3}-6\right)+2}{c\left(3\sqrt{3}\sqrt{\frac{1}{c^{2}}-\frac{4}{c}+3}-9\right)+5}.$ (52)
One of the most remarkable properties of the above equation for the coexistence curve in the reduced parameter space is its independence of the CFT thermodynamic variable $\tilde{\alpha}$, as well as the boundary volume $\mathcal{V}$. This behavior is analogous to the phase transition in the $5$D neutral Gauss-Bonnet AdS black hole, where the coexistence curve is independent of the Gauss-Bonnet coupling constant $\alpha$ when expressed in the reduced parameter space, as discussed in [64]. From the coexistence curve in Fig. 5, the temperature for the low and high entropy coexistence phases decreases with the increasing of the inverse of the central charge $1/C$.
@@ III.3 Critical exponents
$S_{\text{L}}$ $=\sqrt{2}\pi C\sqrt{1-24y-\sqrt{48y(9y-1)+1}}$
×$\times\left(1-12y-\sqrt{48y(9y-1)+1}\right),$ (58)
@@
$\begin{split}\eta&=\frac{S_{\text{H}}-S_{\text{L}}}{S_{\text{c}}}\\ &=\left(12y+\sqrt{48y(9y-1)+1}-1\right)^{2}\\ &\times\frac{36y+\sqrt{48y(9y-1)+1}-1}{8\sqrt{3}y\left(-24y-\sqrt{48y(9y-1)+1}+1\right)^{3/2}}.\end{split}$ (60)
@@
$\begin{split}t&=\frac{\tilde{T}}{\tilde{T}_{\text{c}}}-1\\ &=\sqrt{1-24y-\sqrt{48y(9y-1)+1}}\\ &\times\frac{24y+\sqrt{48y(9y-1)+1}-2}{\sqrt{3}\left(20y+\sqrt{48y(9y-1)+1}-1\right)}-1.\end{split}$ (61)
@@
$\begin{split}\eta&=\frac{4\left(t(t+2)-\sqrt{3}\sqrt{t(t+1)^{2}(t+2)}+1\right)^{2}}{1-2t(t+2)}\\ &\times\frac{\sqrt{3}\sqrt{t(t+1)^{2}(t+2)}-3t(t+2)}{\left(4t(t+2)-2\sqrt{3}\sqrt{t(t+1)^{2}(t+2)}+1\right)^{3/2}}\\ &=4\sqrt{6}t^{\frac{1}{2}}+25\sqrt{6}t^{\frac{3}{2}}+\mathcal{O}(t^{\frac{5}{2}}).\end{split}$ (63)
@@
$\begin{split}\chi\approx\frac{(1+3c)(c-1)^{\frac{1}{3}}}{3c^{\frac{4}{3}}}\sim\left(c-1\right)^{\frac{1}{3}},\end{split}$ (64)
@@
Thus, we have obtained the critical exponents $\alpha^{\prime}$, $\beta$ and $\delta$ associated with the $\tilde{T}-S$ criticality of the CFT, where the conformal factor $\omega$ is free to vary but with fixed Newton’s constant. Our results suggest that these critical exponents associated with the $\tilde{T}-S$ criticality of the CFT are the same as those of the five-dimensional Gauss-Bonnet AdS black hole [42, 65].
IV Conclusion and discussion
In the AdS/CFT correspondence, black holes in AdS spacetime correspond to thermal states of the boundary field theories. Based on the recently proposed viewpoint of the holographic thermodynamics, which treats the conformal factor $\omega$ as a variable, we derived the first law and Euler equation for the thermodynamics of the CFT, and established a holographic dictionary between the thermodynamics of the five-dimensional neutral Gauss-Bonnet AdS black hole and the dual boundary thermodynamics. Having the thermodynamics for the dual CFT, we investigated the phase transition and criticality of the CFT in the fixed $(C,\mathcal{V},\tilde{\alpha})$ canonical ensemble. Just like the free energy for the five-dimensional neutral Gauss-Bonnet AdS black hole, the free energy for the CFT displays a characteristic swallowtail behavior. This indicates a first-order phase transition below the critical point and a second-order phase transition at the critical point. Through the $\tilde{T}-S$ oscillatory behavior, we obtained the critical point for the thermodynamics of the CFT in the $\tilde{T}-S$ plane. The critical point for the CFT is exactly the same as that of the five-dimensional neutral Gauss-Bonnet AdS black hole. Using Maxwell’s equal area law, we got the coexistence curve for the high and low entropy phases of the CFT. Besides, we got the critical exponents for the CFT, and found that these critical exponents associated with the $\tilde{T}-S$ criticality are the same as those of the five-dimensional Gauss-Bonnet AdS black hole. We have chosen the CFT thermodynamical variable $\tilde{\alpha}=1$ in drawing some of our figures, but the qualitative characteristics remain unaffected by the specific values we select.
Critical exponents describe the singularity behavior of thermodynamic variables such as the heat capacity, order parameter, etc. They are universal and satisfy the scaling hypothesis. The critical exponent $\gamma$ governing the isothermal compressibility of black hole thermodynamics is defined as
@@ IV Conclusion and discussion
As we argued in Sec. II.2, the black hole thermodynamic pressure $P$ is dual to the CFT central charge $C$, and the black hole thermodynamic volume $V$ is dual to the CFT chemical potential $\mu$. In our investigation of the critical exponents for the CFT thermodynamics in the $\tilde{T}-S$ plane, we found that if we define the critical exponent $\gamma$ just by replacing the pressure $P$ and the specific volume $v$ by the central charge $C$ and the chemical potential $\mu$ respectively
@@
Criticality is a very important phenomenon for thermodynamic systems, and critical exponents are universal. Our investigation of the thermodynamics and criticality from the gravity/gauge duality viewpoint might lead to new insight into the understanding of the Van der Waals like phase behavior of the five-dimensional Gauss-Bonnet AdS black hole.
We conclude with a final remark on the construction of CFT thermodynamics from bulk thermodynamics for higher-dimensional black holes and modified gravity theories. The number of central charges depends on the spacetime dimension. For higher-dimensional black holes in AdS space, two or more central charges may arise. Additionally, the value of the central charge can vary for different gravitational theories and they might not be the same value. However, the construction method proposed in Ref. [36] yields a single central charge. This observation is worth further investigation.
@@ Acknowledgements.
We are grateful to Yi Pang, Hong Lü, Chuan Chen, Shan-Ping Wu, and Yu-Peng Zhang for useful discussions. We thank the Center for Joint Quantum Studies and Department of Physics (Tianjin University), where part of this work has been done, for its hospitality. This work was supported by National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12305065, No. 12247178, No. 12475056, No. 12347177, No. 12075103 and No. 12247101), the China Postdoctoral Science Foundation (Grant No. 2023M731468), Gansu Province’s Top Leading Talent Support Plan, and the 111 Project (Grant No. B20063).